# Sign changes of Kloosterman sums with moduli having at most two prime factors

### v1

@@ 6. Upper bound for $R_{3}(X)$
The proof of Proposition 2.3 is actually similar to Section 5 in [18]. However, for completeness, we still provide the details of the proof.
@@ 6.1. Separation of variables and establishment of integrals
Recall the Weil’s bound for Kloosterman sums, we get an upper bound of $R_{3}(X)$ which is
$=$ $\sum_{\Pi_{l}=P_{0}P_{1}}\sum_{n\equiv 0(P_{0})\atop(n/P_{0},P_{1})=1}g\left(\frac{n}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{(n,\Pi_{l})=1}g\left(\frac{nP_{0}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{d}\xi_{d}\sum_{n\Pi_{l}\equiv 0(d)\atop(n,\Pi_{l})=1}g\left(\frac{nP_{0}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{d}\xi_{d}\sum_{dk\equiv 0(\Pi_{l})\atop(d,k)=1,(dk/\Pi_{l},\Pi_{l})=1}g\left(\frac{dkP_{0}}{X\Pi_{l}}\right)\mu^{2}\left(\frac{dk}{\Pi_{l}}\right)\kappa^{\omega\left(\frac{dk}{\Pi_{l}}\right)}$
$=$ $\sum_{P_{0}\mid\Pi_{l}}\mu^{2}(P_{0})\kappa^{\omega(P_{0})}\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\xi_{d}$
(5) ×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\sum_{k\equiv 0(\Delta_{2})\atop(k,d)=1,(k/\Delta_{2},\Pi_{l})=1}g\left(\frac{dkP_{0}}{X\Delta_{1}\Delta_{2}}\right)\mu^{2}\left(\frac{dk}{\Delta_{1}\Delta_{2}}\right)\kappa^{\omega\left(\frac{dk}{\Delta_{1}\Delta_{2}}\right)}.$
@@
$\sum_{(k,\Pi_{l})=1,(k,d)=1}g\left(\frac{(d/\Delta_{1})kP_{0}}{X}\right)\mu^{2}\left(\frac{d}{\Delta_{1}}k\right)\kappa^{\omega\left(dk/\Delta_{1}\right)}$
$=$ $\frac{1}{2\pi i}\int_{(2)}\tilde{g}(s)\left(\sum_{(k,(d/\Delta_{1})\Pi_{l})=1}\frac{\mu^{2}(dk/\Delta_{1})\kappa^{\omega(dk/\Delta_{1})}}{(d/\Delta_{1})^{s}k^{s}P_{0}^{s}}\right)X^{s}\text{d}s,$
where the Mellin inverse transform of $\tilde{g}$ is used. Substituting into (6.1) yields
$\frac{X}{2\pi i}\int_{(1)}\tilde{g}(s+1)\left(\sum_{P_{0}\mid\Pi_{l}}\frac{\mu^{2}(P_{0})\kappa^{\omega(P_{0})}}{P_{0}^{s+1}}\right)$ $\left(\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\frac{\xi_{d}\mu^{2}(d/\Delta_{1})\kappa^{\omega(d/\Delta_{1})}}{(d/\Delta_{1})^{s+1}}\right)$
(6) ×$\times\left(\sum_{(k,d\Pi_{l}/\Delta_{1})=1}\frac{\mu^{2}(k)\kappa^{\omega(k)}}{k^{s+1}}\right)X^{s}\text{d}s.$
Furthermore, using the Euler product formula for the Dirichlet series in the third bracket we derive
$\sum_{(k,d\Pi_{l}/\Delta_{1})=1}\frac{\mu^{2}(k)\kappa^{\omega(k)}}{k^{s+1}}$
$=$ $\prod_{p\nmid d\Pi_{l}/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)$
@@
$G(s+1):=\prod_{p}\left(1+\frac{\kappa}{p^{s+1}}\right)\left(1-\frac{1}{p^{s+1}}\right)^{\kappa}.$
The sum in the first bracket of (6.1) is equal to
$\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right),$
@@
$N(s):=\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\frac{\xi_{d}\mu^{2}(d/\Delta_{1})\kappa^{\omega(d/\Delta_{1})}}{(d/\Delta_{1})^{s+1}}\prod_{p\mid d/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
@@ 6.2. Analysis of $N(s)$
Obviously, there is
$N(s)=\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{(d,\Pi_{l})=1}\frac{\xi_{d\Delta_{1}}\mu^{2}(d)\kappa^{\omega(d)}}{d^{s+1}}\prod_{p\mid d}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
We define
$\beta(d,s):=\frac{d^{s+1}}{\mu^{2}(d)\kappa^{\omega(d)}}\prod_{p\mid d}\left(1+\frac{\kappa}{p^{s+1}}\right),$
then from the definition of $\xi_{d}$ we have
$N(s)=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{(d,\Pi_{l})=1}\frac{\xi_{d\Delta_{1}}}{\beta(d,s)}$
$=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{(d,\Pi_{l})=1}\left(\sum_{d\Delta_{1}=[d_{1},d_{2}]}\frac{\lambda_{d_{1}}\lambda_{d_{2}}}{\beta(d,s)}\right)$
$=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{[d_{1},d_{2}]\equiv 0(\Delta_{1})\atop([d_{1},d_{2}]/\Delta_{1},\Pi_{l})=1}\frac{\lambda_{d_{1}}\lambda_{d_{2}}}{\beta([d_{1},d_{2}]/\Delta_{1},s)}.$
Since $d_{i}(i=1,2)$ are square-free integers, let $(d_{1},d_{2})=m$, $d_{1}=d_{1}^{\prime}m$ and $d_{2}=d_{2}^{\prime}m$. Then we know $m$, $d_{1}^{\prime}$ and $d_{2}^{\prime}$ are pairwise coprime (for convenience, we will use the symbol $d_{1}$, $d_{2}$ in the following text). Now we can write
$N(s)=\sum_{\Delta_{1}\mid\Pi_{l}}\sum_{m\leq\sqrt{D}}\sum_{md_{1}d_{2}\equiv 0(\Delta_{1})\atop(d_{1},d_{2})=1,(md_{1}d_{2}/\Delta_{1},\Pi_{l})=1}\frac{\lambda_{md_{1}}\lambda_{md_{2}}}{\beta(md_{1}d_{2}/\Delta_{1},s)}.$
@@
$N(s)=$ $\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\sum_{m\leq\sqrt{D}}\sum_{m\equiv 0(\Delta_{1})\atop{d_{1}\equiv 0(\Delta_{2})\atop{d_{2}\equiv 0(\Delta_{3})\atop(d_{1},d_{2})=1,\left(\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\frac{d_{2}}{\Delta_{3}},\Pi_{l}\right)=1}}}\frac{\lambda_{md_{1}}\lambda_{md_{2}}}{\beta\left(\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\frac{d_{2}}{\Delta_{3}},s\right)}$
(8) $=$ $\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{1}{\beta(m/\Delta_{1},s)}\sum_{d_{1}\equiv 0(\Delta_{2})\atop(d_{1}/\Delta_{2},\Pi_{l})=1}\frac{\lambda_{md_{1}}}{\beta(d_{1}/\Delta_{2},s)}\sum_{d_{2}\equiv 0(\Delta_{3})\atop{(d_{1},d_{2})=1\atop(d_{2}/\Delta_{3},\Pi_{l})=1}}\frac{\lambda_{md_{2}}}{\beta(d_{2}/\Delta_{3},s)}.$
Firstly, we deal with the inner sum in (6.2). From the definition of $\lambda_{d}$ we have
$\sum_{d_{2}\equiv 0(\Delta_{3})\atop{(d_{1},d_{2})=1\atop(d_{2}/\Delta_{3},\Pi_{l})=1}}\frac{\lambda_{md_{2}}}{\beta(d_{2}/\Delta_{3},s)}$
$=$ $\sum_{(d_{2},\Pi_{l})=1\atop(d_{1},d_{2})=1}\frac{\lambda_{md_{2}\Delta_{3}}}{\beta(d_{2},s)}$
$=$ $\mu(m\Delta_{3})\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)}F\left(\frac{\log(\sqrt{D}/md_{2}\Delta_{3})}{\log\sqrt{D}}\right).$
Set
$P(x):=F\left(x\cdot\frac{\log(\sqrt{D}/m\Delta_{3})}{\log\sqrt{D}}\right),$
and apply Lemma 3.3 to $P(x)$ (or see Page 1218 in [18]), we obtain
$\frac{\mu(m\Delta_{3})}{2\pi i}\int_{(2)}\check{P}_{\sqrt{D}/m\Delta_{3}}(t)\left(\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)d_{2}^{t}}\right)\frac{(\sqrt{D}/(m\Delta_{3}))^{t}}{t}\text{d}t.$
Note that
$\check{P}_{\sqrt{D}/m\Delta_{3}}(t)=\check{F}_{\sqrt{D}}(t),$
@@
(9) $=$ $\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}\prod_{p\mid(d_{1}/\Delta_{2})}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}.$
Substituting into (6.2) to derive
$\frac{1}{2\pi i}\int_{(2)}$ $\check{F}_{\sqrt{D}}(t_{1})\frac{\sqrt{D}^{t_{1}}}{t_{1}}\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)\right)$
@@
Next, following much the same way, we deal with the inner sum in (6.2), which leads to
@@
Thus we obtain that (6.2) is further equal to
$\frac{1}{(2\pi i)^{2}}\iint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}}}{t_{1}+t_{2}}$
×$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}\Delta_{2}^{t_{2}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)\right)$
×$\times\left(\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{\mu^{2}(m)}{\beta(m/\Delta_{1},s)m^{t_{1}+t_{2}}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right)\text{d}t_{1}\text{d}t_{2}$
$=\frac{1}{(2\pi i)^{3}}\iiint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}$
×$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}\Delta_{2}^{t_{2}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)\right)$
×$\times\left(\sum_{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}\frac{\mu^{2}(m)}{\beta(m/\Delta_{1},s)m^{t_{1}+t_{2}+w}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w,$
where the third integral comes from the Mellin inverse transform of $1_{m\leq\sqrt{D}}$.
@@
which leads to
$N(s)=\frac{1}{(2\pi i)^{3}}\iiint$ $\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}$
×$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu^{2}(\Delta_{1})\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{1}^{t_{1}+t_{2}+w}\Delta_{2}^{t_{2}}\Delta_{3}^{t_{1}}}\right)$
×$\times\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w.$

### v2

@@ 5. Proof of the first Proposition
The estimation of $T(X)$ is actually similar to Section 5 in [17]. However, for completeness, we provide the details of the proof below.
@@ 5.1. Separation of variables and establishment of integrals
We write $T(X)$ as
$\sum_{(n,\Pi_{l})=1}g\left(\frac{n\Pi_{l}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}\left(\sum_{d\mid n\Pi_{l}}\lambda_{d}\right)^{2}$
$=$ $\sum_{d}\xi_{d}\sum_{n\Pi_{l}\equiv 0(d)\atop(n,\Pi_{l})=1}g\left(\frac{n\Pi_{l}}{X}\right)\mu^{2}(n)\kappa^{\omega(n)}$
$=$ $\sum_{d}\xi_{d}\sum_{dk\equiv 0(\Pi_{l})\atop(d,k)=1,(dk/\Pi_{l},\Pi_{l})=1}g\left(\frac{dk}{X}\right)\mu^{2}\left(\frac{dk}{\Pi_{l}}\right)\kappa^{\omega\left(\frac{dk}{\Pi_{l}}\right)}$
$=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\xi_{d}\sum_{k\equiv 0(\Delta_{2})\atop(k,d)=1,(k/\Delta_{2},\Pi_{l})=1}g\left(\frac{dk}{X}\right)\mu^{2}\left(\frac{dk}{\Delta_{1}\Delta_{2}}\right)\kappa^{\omega\left(\frac{dk}{\Delta_{1}\Delta_{2}}\right)}$
(11) $=$ $\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\xi_{d}\sum_{(k,d)=1\atop(k,\Pi_{l})=1}g\left(\frac{dk\Delta_{2}}{X}\right)\mu^{2}\left(\frac{d}{\Delta_{1}}k\right)\kappa^{\omega\left(dk/\Delta_{1}\right)}.$
@@
$\frac{1}{2\pi i}\int_{(2)}\tilde{g}(s)\left(\sum_{(k,(d/\Delta_{1})\Pi_{l})=1}\frac{\mu^{2}(dk/\Delta_{1})\kappa^{\omega(dk/\Delta_{1})}}{d^{s}k^{s}\Delta_{2}^{s}}\right)X^{s}\text{d}s,$
where the Mellin inverse transform of $\tilde{g}$ is used. Substituting the above into (5.1) yields
(12) $\frac{X}{2\pi i}\int_{(1)}\tilde{g}(s+1)\left(\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\frac{1}{\Delta_{2}^{s+1}}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\frac{\xi_{d}\mu^{2}(d/\Delta_{1})\kappa^{\omega(d/\Delta_{1})}}{d^{s+1}}\right)\left(\sum_{(k,d\Pi_{l}/\Delta_{1})=1}\frac{\mu^{2}(k)\kappa^{\omega(k)}}{k^{s+1}}\right)X^{s}\text{d}s.$
Furthermore, using the Euler product formula for the Dirichlet series in the third bracket, for $\Re s>0$, we derive
$\sum_{(k,d\Pi_{l}/\Delta_{1})=1}\frac{\mu^{2}(k)\kappa^{\omega(k)}}{k^{s+1}}=\prod_{p\nmid d\Pi_{l}/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)$
@@
$G(s+1):=\prod_{p}\left(1+\frac{\kappa}{p^{s+1}}\right)\left(1-\frac{1}{p^{s+1}}\right)^{\kappa},$
@@
$N(s):=\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\frac{1}{\Delta_{2}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\sum_{d\equiv 0(\Delta_{1})\atop(d/\Delta_{1},\Pi_{l})=1}\frac{\xi_{d}\mu^{2}(d/\Delta_{1})\kappa^{\omega(d/\Delta_{1})}}{d^{s+1}}\prod_{p\mid d/\Delta_{1}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
@@ 5.2. Analysis of $N(s)$
Obviously, we have
$N(s)=\sum_{\Pi_{l}=\Delta_{1}\Delta_{2}}\frac{1}{\Delta_{1}^{s+1}\Delta_{2}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\sum_{(d,\Pi_{l})=1}\frac{\xi_{d\Delta_{1}}\mu^{2}(d)\kappa^{\omega(d)}}{d^{s+1}}\prod_{p\mid d}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}.$
Define
$\beta(d,s):=\frac{d^{s+1}}{\mu^{2}(d)\kappa^{\omega(d)}}\prod_{p\mid d}\left(1+\frac{\kappa}{p^{s+1}}\right).$
Then from the definition of $\xi_{d}$, we get
$\sum_{(d,\Pi_{l})=1}\frac{\xi_{d\Delta_{1}}}{\beta(d,s)}=\sum_{(d,\Pi_{l})=1}\left(\sum_{d\Delta_{1}=[d_{1},d_{2}]}\frac{\lambda_{d_{1}}\lambda_{d_{2}}}{\beta(d,s)}\right)=\sum_{[d_{1},d_{2}]\equiv 0(\Delta_{1})\atop([d_{1},d_{2}]/\Delta_{1},\Pi_{l})=1}\frac{\lambda_{d_{1}}\lambda_{d_{2}}}{\beta([d_{1},d_{2}]/\Delta_{1},s)}.$
Since $d_{i}(i=1,2)$ are square-free integers, let $(d_{1},d_{2})=m$, $d_{1}=d_{1}^{\prime}m$ and $d_{2}=d_{2}^{\prime}m$; then we know $m$, $d_{1}^{\prime}$ and $d_{2}^{\prime}$ are pairwise coprime (for convenience, we will still use the symbol $d_{1}$, $d_{2}$ in the following text). Now we can write
$N(s)=\sum_{\Delta_{1}\mid\Pi_{l}}\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\sum_{m\leq\sqrt{D}}\sum_{md_{1}d_{2}\equiv 0(\Delta_{1})\atop(d_{1},d_{2})=1,(md_{1}d_{2}/\Delta_{1},\Pi_{l})=1}\frac{\lambda_{md_{1}}\lambda_{md_{2}}}{\beta(md_{1}d_{2}/\Delta_{1},s)}.$
@@
$N(s)=$ $\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\sum_{m\leq\sqrt{D}}\sum_{m\equiv 0(\Delta_{1})\atop{d_{1}\equiv 0(\Delta_{2})\atop{d_{2}\equiv 0(\Delta_{3})\atop(d_{1},d_{2})=1,\left(\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\frac{d_{2}}{\Delta_{3}},\Pi_{l}\right)=1}}}\frac{\lambda_{md_{1}}\lambda_{md_{2}}}{\beta\left(\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\frac{d_{2}}{\Delta_{3}},s\right)}$
$=$ $\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{1}{\beta(m/\Delta_{1},s)}$
(14) $\times\sum_{d_{1}\equiv 0(\Delta_{2})\atop(d_{1}/\Delta_{2},\Pi_{l})=1}\frac{\lambda_{md_{1}}}{\beta(d_{1}/\Delta_{2},s)}\sum_{d_{2}\equiv 0(\Delta_{3})\atop{(d_{1},d_{2})=1\atop(d_{2}/\Delta_{3},\Pi_{l})=1}}\frac{\lambda_{md_{2}}}{\beta(d_{2}/\Delta_{3},s)}.$
Firstly, we deal with the inner sum in (5.2). From the definition of $\lambda_{d}$, we have
$\sum_{d_{2}\equiv 0(\Delta_{3})\atop{(d_{1},d_{2})=1\atop(d_{2}/\Delta_{3},\Pi_{l})=1}}\frac{\lambda_{md_{2}}}{\beta(d_{2}/\Delta_{3},s)}=\sum_{(d_{1},d_{2})=1\atop(d_{2},\Pi_{l})=1}\frac{\lambda_{md_{2}\Delta_{3}}}{\beta(d_{2},s)}$
$=$ $\mu(m\Delta_{3})\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)}F\left(\frac{\log(\sqrt{D}/(md_{2}\Delta_{3}))}{\log\sqrt{D}}\right).$
We replace variable $d_{2}$ with $d_{2}\Delta_{3}$ (we still use the symbol $d_{2}$) here. From (5.2), one finds the restrictions $d_{2}\equiv 0(\Delta_{3})$ and $(d_{1},d_{2})=1$, which yields $(d_{1},\Delta_{3})=1$. We released condition $(d_{1},\Delta_{3})=1$, which is included in condition $(d_{1}/\Delta_{2},\Pi_{l})=1$ because the numbers discussed here are all square-free numbers. The same operations in the following text will not be explained again.
Setting
$S(x):=F\left(x\cdot\frac{\log(\sqrt{D}/(m\Delta_{3}))}{\log\sqrt{D}}\right),$
and applying Lemma 3.6 to $S(x)$ (or see Page 1218 in [17]), we obtain
$\frac{\mu(m\Delta_{3})}{2\pi i}\int_{(2)}\check{S}_{\sqrt{D}/(m\Delta_{3})}(t)\left(\sum_{\left(d_{2},\frac{m}{\Delta_{1}}\frac{d_{1}}{\Delta_{2}}\Pi_{l}\right)=1}\frac{\mu(d_{2})}{\beta(d_{2},s)d_{2}^{t}}\right)\frac{(\sqrt{D}/(m\Delta_{3}))^{t}}{t}\text{d}t.$
Noting that
$\check{S}_{\sqrt{D}/(m\Delta_{3})}(t)=\check{F}_{\sqrt{D}}(t),$
@@
$=$ $\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}\prod_{p\mid(d_{1}/\Delta_{2})}\left(1-\frac{1}{\beta(p,s)p^{t}}\right)^{-1}.$
Substituting the above into (5.2) gives
$\frac{1}{2\pi i}\int_{(2)}$ $\check{F}_{\sqrt{D}}(t_{1})\frac{\sqrt{D}^{t_{1}}}{t_{1}}\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}}{\Pi_{l}^{s+1}}\frac{\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}\right)\right)$
@@
Next, in much the same way, we handle the inner sum in (5.2), which yields
@@
Thus we obtain that (5.2) is equal to
$\frac{1}{(2\pi i)^{2}}\iint_{(2)}\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}}}{t_{1}+t_{2}}\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}$
$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}\Delta_{2}^{t_{2}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)\right)$
$\times\left(\sum_{m\leq\sqrt{D}\atop{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}}\frac{\mu^{2}(m)}{\beta(m/\Delta_{1},s)m^{t_{1}+t_{2}}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right)\text{d}t_{1}\text{d}t_{2}$
$=$ $\frac{1}{(2\pi i)^{3}}\iiint_{(2)}\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}\frac{1}{\Pi_{l}^{s+1}}\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}$
$\times\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{3}^{t_{1}}\Delta_{2}^{t_{2}}}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)\right)$
$\times\left(\sum_{m\equiv 0(\Delta_{1})\atop(m/\Delta_{1},\Pi_{l})=1}\frac{\mu^{2}(m)}{\beta(m/\Delta_{1},s)m^{t_{1}+t_{2}+w}}\prod_{p\mid(m/\Delta_{1})}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}\right)^{-1}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w,$
where $\iint_{(2)}$, $\iiint_{(2)}$ represent $\int_{(2)}\int_{(2)}$ and $\int_{(2)}\int_{(2)}\int_{(2)}$, respectively. The third integral comes from the Mellin inverse transform of
$\mathbbm{1}_{m\leq\sqrt{D}}=\begin{cases}&1,\ \text{if }m\leq\sqrt{D},\\ &0,\ \text{if }m>\sqrt{D}.\end{cases}$
@@
and this leads to
$N(s)=$ $\frac{1}{(2\pi i)^{3}}\iiint_{(2)}\check{F}_{\sqrt{D}}(t_{1})\check{F}_{\sqrt{D}}(t_{2})\frac{\sqrt{D}^{t_{1}+t_{2}+w}}{t_{1}+t_{2}+w}\frac{1}{\Pi_{l}^{s+1}}\left(\sum_{\Delta_{1}\Delta_{2}\Delta_{3}\mid\Pi_{l}}\frac{\mu^{2}(\Delta_{1})\mu(\Delta_{2})\mu(\Delta_{3})}{\Delta_{1}^{t_{1}+t_{2}+w}\Delta_{2}^{t_{2}}\Delta_{3}^{t_{1}}}\right)$
$\times\prod_{p\mid\Pi_{l}}\left(1+\frac{\kappa}{p^{s+1}}\right)^{-1}\prod_{p\nmid\Pi_{l}}\left(1-\frac{1}{\beta(p,s)p^{t_{1}}}-\frac{1}{\beta(p,s)p^{t_{2}}}+\frac{1}{\beta(p,s)p^{t_{1}+t_{2}+w}}\right)\text{d}t_{1}\text{d}t_{2}\text{d}w.$